OpenStudy (emogirlatemoooocow):

A chef boiled 4 eggs and put them in a basket with 8 eggs that were not boiled. All of the eggs look the same. Randy selects an egg, keeps it, and then selects another egg. Which expression gives the probability that he selects 2 boiled eggs?

OpenStudy (emogirlatemoooocow):

A. \[\frac{ 4 }{ 12 } • \frac{ 4 }{ 12}\]

OpenStudy (emogirlatemoooocow):

B.\[\frac{ 4 }{ 12 } • \frac{ 3 }{ 12 }\]

OpenStudy (emogirlatemoooocow):

C.\[\frac{ 4 }{ 12 } • \frac{ 3 }{ 11 }\]

OpenStudy (myth1111):

Well, how many eggs were there in total before Randy took one

OpenStudy (emogirlatemoooocow):

D.\[\frac{ 4 }{ 12 } • \frac{ 4 }{ 11 }\]

OpenStudy (emogirlatemoooocow):

12 @myth1111

OpenStudy (mattga2002):

16/132

OpenStudy (myth1111):

Right don't listen to that guy lol

OpenStudy (myth1111):

Right so there was 12 eggs in total, and 4 boiled ones, so the first fraction is....

OpenStudy (mattga2002):

Remember to reduce

OpenStudy (emogirlatemoooocow):

\[\frac{ 4 }{ 12 } or \frac{ 12 }{ 4 } ?\]

OpenStudy (myth1111):

Well, which makes more sense to you?

OpenStudy (mattga2002):

The answer is 4/12

OpenStudy (emogirlatemoooocow):

im thinking 4 over 12

OpenStudy (myth1111):

Yeah, it's 4/12. Mattga you are not allowed to just give answers. Also you are wrong.

OpenStudy (mattga2002):

It is because he cannot have a probability over 1/1

OpenStudy (emogirlatemoooocow):

@mattga2002 pls stop :|

OpenStudy (emogirlatemoooocow):

your gonna get yourself banned .,.

OpenStudy (myth1111):

So now, for the 2nd fraction... there was 12 eggs, and 1 was taken, so how many eggs are there now?

OpenStudy (emogirlatemoooocow):

11

OpenStudy (emogirlatemoooocow):

i think

OpenStudy (myth1111):

Yep

OpenStudy (myth1111):

And since there was 4 boiled eggs, and he took one, how many boiled eggs are there?

OpenStudy (emogirlatemoooocow):

3

OpenStudy (myth1111):

Yup, so 3, and 11, so what's the 2nd fraction?

OpenStudy (emogirlatemoooocow):

\[\frac{ 3 }{ 11}\]

OpenStudy (myth1111):

So the answer is?

OpenStudy (emogirlatemoooocow):

Or 11| 3

OpenStudy (myth1111):

it's 3/11

OpenStudy (emogirlatemoooocow):

\[\frac{ 11 }{ 3 }\]

OpenStudy (emogirlatemoooocow):

Thanks :) everyone who helped <3

OpenStudy (myth1111):

btw in these probability problems the fractions shouldn't be over 1/1 ever.